OpenStudy (anonymous):

Part 1: Explain, in complete sentences, how you would use the elimination method to solve the following system of equations. Part 2: Provide the solution to the system. 2x – 5y = –12 3x – 4y = 3

OpenStudy (anonymous):

subtract line 1 from line two which will give you a singular x and y (x+y=15) then solve for either x or y and substitute into one of the original equations. x=15-y (30-3y)-5y=-12 -8y=-42